# Qubit encoding for a mixture of localized functions

### v1

@@ Abstract
One of the crucial generic techniques for quantum computation is the amplitude encoding. Although such techniques have been proposed, each of them often requires exponential classical-computational cost or an oracle whose explicit construction is not provided. Given the recent demands for practical quantum computation, we develop moderately specialized encoding techniques that generate an arbitrary linear combination of localized complex functions. We demonstrate that $n_{\mathrm{loc}}$ discrete Lorentzian functions as an expansion basis set lead to efficient probabilistic encoding, whose computational time is $\mathcal{O}(\max(n_{\mathrm{loc}}^{2}\log n_{\mathrm{loc}},n_{\mathrm{loc}}^{2}\log n_{q},n_{q}\log n_{q}))$ for $n_{q}$ data qubits equipped with $\log_{2}n_{\mathrm{loc}}$ ancillae. Furthermore, amplitude amplification in combination with amplitude reduction renders it deterministic with controllable errors and the computational time is reduced to $\mathcal{O}(\max(n_{\mathrm{loc}}^{3/2}\log n_{\mathrm{loc}},n_{\mathrm{loc}}^{3/2}\log n_{q},n_{q}\log n_{q})).$ We provide estimation of required resources for application of our scheme to quantum chemistry in real space. We also show the results on real superconducting quantum computers to confirm the validity of our techniques.
@@ I Introduction
We often encounter a case where a quantum algorithm that is mathematically proven to solve a problem more efficiently than classical computers assumes that an initial many-qubit state has already been prepared in which the initial condition is appropriately encoded in some format specified by the algorithm. Such generic techniques for generating an arbitrary many-qubit state from initialized qubits as efficiently as possible are called the amplitude encoding. Since an $n$-qubit system has approximately $2^{n}$ degrees of freedom, the encoding of a truly arbitrary state, i.e., a state under no constraints, using predetermined circuit parameters inevitably suffers from exponential classical-computational cost with respect to $n$ [1, 2, 3, 4, 5, 6]. If we exploit quantum parallelism, on the other hand, exponential cost can be avoided. There are efficient encoding techniques for arbitrary states under no or only weak constraints [7, 8, 9]. Each of them assumes, however, that an oracle capable of performing an elaborate unitary operation on $\mathcal{O}(2^{n})$ independent states simultaneously is available. This fact prohibits us from finding an efficient encoding scheme of practical use of today.
@@
When we tackle a specific kind of problems, however, such techniques for arbitrary states are often found to be too generic, that is, plausible initial states for the problems are characterized by much less number of parameters than $2^{n}.$ Let us consider, for example, a quantum algorithm that admits an initial state in which a probability distribution function is encoded. Even if we want to try various normal distributions as the inputs, a generic encoding technique demands exponential cost, despite the states being specified only by their means and widths. Intuitively speaking, the degrees of freedom in a generic encoding technique is unnecessarily enormous compared to the amount of information specifying an initial state of practical use. Motivated by this consideration, we develop moderately specialized encoding techniques in the present study: generation of an arbitrary linear combination (LC) of localized complex functions.
@@
To this end, we design the main encoding technique by starting from the probabilistic operation [10, 11] for an LC of discrete Lorentzian functions (LFs), to which we apply the quantum amplitude amplification (QAA) technique [12, 13] to render the encoding deterministic. As is demonstrated later, the expansion of a target function in LFs is favorable for achieving efficient circuit implementation. Our encoding techniques possess applicability to diverse fields of quantum computation. Amongst them, quantum chemistry in real space [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24] is a promising one, where an initial state can be constructed from one-electron molecular orbitals (MOs) [25, 26], typically expressed as LCs of localized orbitals. We provide resource estimation for such calculations. To demonstrate the efficiency and practical usefulness of our scheme, we perform it on the real quantum computers commercially provided by IBM. Our encoding scheme is for a dense quantum state, that is, it treats a target function that has basically a nonzero amplitude of each computational basis. For encoding schemes for a sparse quantum state, see, e.g., Refs [27, 28, 29] and references therein.
@@
It is noted that there exist the encoding techniques based on training of parametrized circuits [30, 31]. Although such techniques are promising for practical use, we focus on those not based on training in the present study. Also, we note here that Moosa et al.[32] proposed the Fourier series loader method. It encodes first the Fourier coefficients of a target function on a coarse mesh, after which they are reinterpreted as those for a finer mesh to be Fourier transformed backward. See also Ref. [33] for the quantum interpolation method.
@@ II.1.1 Probabilistic operation of an LC of unitaries and amplitude reduction
Figure 1: Each of circuits (a) $\mathcal{C}_{\mathrm{prob}}$ and (b) $\mathcal{C}_{\mathrm{prob}}^{(\mathrm{AR})}$ is for probabilistic operation of an LC $\mathcal{A}$ of unitaries. The explicit construction of the central part $\mathcal{C}$ in $\mathcal{C}_{\mathrm{prob}}$ is described in the original paper [11]. $\mathcal{C}_{\mathrm{prob}}^{(\mathrm{AR})}$ implements the amplitude reduction for the success state by introducing an extra ancilla in addition to those in $\mathcal{C}_{\mathrm{prob}}.$
@@ II.1.2 Determinization via amplitude reduction and amplification
The generic technique of QAA [12, 13] amplifies the weight of ‘good’ states contained in an input state by rotating it within the two-dimensional subspace spanned by the good states and the orthogonal states, leading to the quadratic speedup for obtaining the good state. This rotation is achieved by the appropriate number of applications of the amplification operator to the input state. In the present study, we adopt this technique for $|\Psi_{\mathrm{AR}}\rangle$ in $\mathcal{C}_{\mathrm{prob}}^{(\mathrm{AR})}$ to raise the success probability. The good states in this case are all those in which the ancillary part is $|0\rangle_{n_{\mathrm{A}}+1}.$ The amplification operator $Q$ for this case can be implemented according to the generic theory. We assume that we already know the weight $w$ of the desired state $|\psi_{\mathrm{lc}}\rangle$ in $\mathcal{C}_{\mathrm{prob}}$ precisely by employing techniques of quantum amplitude estimation (QAE)[13, 34, 35, 36, 37, 38, 39]. Let us consider $m$ applications of $Q$ to $|\Psi_{\mathrm{AR}}\rangle$ by appending the amplification circuit to $\mathcal{C}_{\mathrm{prob}}^{(\mathrm{AR})}.$ By defining $\theta_{w}$ via $w=\sin^{2}\theta_{w}\ (0<\theta_{w}<\pi/2),$ the weight of good state after the amplification is written as $w_{m}(\theta_{w})\equiv\sin^{2}((2m+1)\theta_{w})$ [12, 13]. If $w_{m}(\theta_{w})$ is equal to unity, the desired state will be obtained with certainty, that is, the nonunitary operation of $\mathcal{A}$ is now deterministic. The operation is still probabilistic otherwise. Specifically, for a fixed $m,$ the following $m$ values of $\theta_{w},$
@@
These observations indicate that $\mathcal{C}_{\mathrm{prob}}^{(\mathrm{AR})}$ with $\theta_{\mathrm{AR,opt}}$ and the subsequent QAA render the nonunitary operation of $\mathcal{A}$ deterministic regardless of $w.$ We refer to this treatment as the determinization of the probabilistic operation and denote it by the quantum amplitude reduction and amplification (QARA) technique. Since $w$ is on the order of $1/n_{\mathrm{loc}},$ we have $m_{\mathrm{opt}}=\mathcal{O}(\sqrt{n_{\mathrm{loc}}})$ from Eq. (4). In particular, for a case where $1/4\leq w<1,$ we have $m_{\mathrm{opt}}=1.$ Nishi et al. [40] have proposed a similar technique for determinizing the QAA for probabilistic imaginary-time evolution (PITE)[18, 20, 41, 22] steps.
@@ II.1.3 Error analysis for inaccurate QAE
Although we have adopted QARA for the determinization, the weight $w$ of a success state can also be raised to unity by employing the technique in the original paper [13] of QAA, which modifies the reflection circuit at the final iteration. We denote this technique by the modified QAA. Thereby a complex transcendental equation has to be solved by using $w$ obtained via QAE for deciding the circuit parameters, and the dependence of the error in the amplified state on that in the QAE process is thus difficult to evaluate. Also, the fixed-point search (FPS) technique [42] amplifies the unknown weight $w$ of a success state to $1-\delta^{2}$ for a specified tolerance $\delta.$ $\delta$ must be strictly larger than zero for achieving the quadratic speedup. The modified QAA and FPS techniques do not require an additional qubit in contrast to QARA.
Let us consider a practical situation where the observer obtains an erroneous estimation $w^{(\mathrm{est})}=w+\Delta w$ of the weight $w$ of the original success state. The circuit parameter $\theta_{\mathrm{AR,opt}}$ for amplitude reduction and the repetition number $m_{\mathrm{opt}}$ for amplitude amplification can now involve errors. We demonstrate in Appendix A that QARA allows us to derive a simple relation between the error in the QAE process and that in the determinization. When we approach the limit of $w\to 0$ with keeping $\Delta w/w$ finite, an approximate maximum of the weight of the failure state after the completion of the QARA process is given by
@@
Figure 2 shows the weight $W_{\mathrm{f}}$ of failure state after the QARA process for $\Delta w/w=0.1,0.04,$ and $0.01$ as a function of $w.$ $W_{\mathrm{f}}$ after the QAA process (without amplitude reduction) is also shown in the figure. While the values of $W_{\mathrm{f}}$ from QAA are seen to scatter for each of $\Delta w/w,$ those from QARA exhibit much smaller variations over the wide range of $w.$ Furthermore, $\epsilon_{\mathrm{QARA}}$ approximates $W_{\mathrm{f}}$ well. These observations suggest that QARA works as an error-controllable method for determinizing a probabilistic operation.
Figure 2: Each circle represents the weight $W_{\mathrm{f}}$ of failure state in the resultant state of QARA with a QAE error $\Delta w$ as a function of the true weight $w$ of success state in the original state. Each cross represents $W_{\mathrm{f}}$ for QAA with a QAE error. Horizontal lines represent the approximate maximum errors in Eq. (6) for the values of $\Delta w/w.$
@@ II.2.2 Displaced single basis function
is easily confirmed to perform modular addition for a computational basis as $T(k)|j\rangle_{n_{q}}=|(j+k)\ \mathrm{mod}\ N\rangle_{n_{q}}.$ The displaced basis function can thus be generated by locating the function at the origin and translate it by $k_{\mathrm{c}\ell}:|\widetilde{f}_{\ell}\rangle=T(k_{\mathrm{c}\ell})U_{\mathrm{orig}}^{(\ell)}|0\rangle_{n_{q}},$ as depicted in Fig. 3(a).
@@ II.2.3 LC of displaced basis functions
with $U_{\mathrm{shift}}^{(\ell)}\equiv U_{\mathrm{shift}}(k_{\mathrm{c}\ell})$ for each $\ell.$ With the LC
$\mathcal{L}\equiv\sum_{\ell=0}^{n_{\mathrm{loc}}-1}d_{\ell}\widetilde{U}^{(\ell)}$ (12)
of them, the desired state in Eq. (7) is written as
$|\psi_{\mathrm{lc}}\rangle=\sum_{\ell=0}^{n_{\mathrm{loc}}-1}d_{\ell}|\widetilde{f}_{\ell}\rangle=\mathrm{QFT}\cdot\mathcal{L}|0\rangle_{n_{q}}.$ (13)
$\mathcal{L},$ which is nonunitary in general, admits the probabilistic implementation as a special case of $\mathcal{C}_{\mathrm{prob}}$ in Fig. 1(a). Specifically, the circuit $\mathcal{C}_{\mathrm{lc}}$ in Fig. 3(b) for the data register and
@@
ancillae encodes $|\psi_{\mathrm{lc}}\rangle$ probabilistically. The central part, $\mathcal{C}_{\mathcal{L}},$ implements $\mathcal{L}$ probabilistically by recognizing the $n_{\mathrm{A}}$-bit integer $\ell$ specified by an ancillary state $|\ell\rangle_{n_{\mathrm{A}}}.$ For details, see the original paper [11], that describes the implementation of a generic LC of unitaries. It is noted here that a slight modification to the circuit in the original paper exists in $\mathcal{C}_{\mathcal{L}}.$ That is, since $\mathrm{QFT}^{\dagger}$ in $\widetilde{U}^{(\ell)}$ does not refer to $\ell$ [see Eq. (11)], the controlled $\mathrm{QFT}^{\dagger}$ gates have been replaced by the uncontrolled $\mathrm{QFT}^{\dagger}$ gate, as seen in Fig. 3(b). Since the $\mathrm{QFT}$ operation after $\mathcal{C}_{\mathcal{L}}$ does not act on the ancillae, it is also possible to perform the measurement, observe the success state, and then operate $\mathrm{QFT}$ on the data register.
Let us estimate the depth of $\mathcal{C}_{\mathrm{lc}}$ by assuming that of $U_{\mathrm{orig}}^{(\ell)}$ is independent of $\ell.$ Since $\mathrm{depth}(\mathrm{QFT})=\mathcal{O}(n_{q}\log n_{q})$ (see Appendix C) and $\mathrm{depth}(U_{\mathrm{shift}}^{(\ell)})=\mathcal{O}(1)$ for each $\ell,$ we find
$\mathrm{depth}\left(\mathcal{C}_{\mathrm{lc}}\right)$ $=\mathcal{O}\left(\max\left(n_{\mathrm{loc}}\mathrm{depth}\left(\mathrm{C}^{n_{\mathrm{A}}}U_{\mathrm{orig}}^{(\ell)}\right),n_{q}\log n_{q}\right)\right).$ (15)
Although we have assumed the basis functions and the coefficients to be real, it is possible to encode any LC of complex basis functions via complex coefficients by redefining the unitaries and coefficients, as explained in Appendix D. The descriptions below for real values can thus be applicable to complex values with small modifications for $n_{\mathrm{A}}+1$ ancillae. Also, the protocol described above can be extended to multidimensional space by employing product basis functions, as explained in Appendix E.

### v2

@@ Abstract
One of the crucial generic techniques for quantum computation is amplitude encoding. Although several approaches have been proposed, each of them often requires exponential classical-computational cost or an oracle whose explicit construction is not provided. Given the growing demands for practical quantum computation, we develop moderately specialized encoding techniques that generate an arbitrary linear combination of localized complex functions. We demonstrate that $n_{\mathrm{loc}}$ discrete Lorentzian functions as an expansion basis set lead to efficient probabilistic encoding, whose computational time is $\mathcal{O}(\max(n_{\mathrm{loc}}^{2}\log n_{\mathrm{loc}},n_{\mathrm{loc}}^{2}\log n_{q},n_{q}))$ for $n_{q}$ data qubits equipped with $\log_{2}n_{\mathrm{loc}}$ ancillae. Furthermore, amplitude amplification in combination with amplitude reduction renders it deterministic analytically with controllable errors and the computational time is reduced to $\mathcal{O}(\max(n_{\mathrm{loc}}^{3/2}\log n_{\mathrm{loc}},n_{\mathrm{loc}}^{3/2}\log n_{q},n_{q})).$ We estimate required resources for applying our scheme to quantum chemistry in real space. We also show the results on real superconducting quantum computers to confirm the validity of our techniques.
@@ I Introduction
We often encounter a case where a quantum algorithm that is mathematically proven to solve a problem more efficiently than classical algorithms assumes that an initial many-qubit state has already been prepared in which the initial condition is appropriately encoded in some format specified by the algorithm. Such generic techniques for preparing an arbitrary many-qubit state from initialized qubits as efficiently as possible are called amplitude encoding. Since an $n$-qubit system has approximately $2^{n}$ degrees of freedom, encoding a truly arbitrary state, i.e., a state under no constraints, using predetermined circuit parameters inevitably suffers from exponential classical-computational cost with respect to $n$ [1, 2, 3, 4, 5, 6]. If we exploit quantum parallelism, on the other hand, exponential cost can be avoided. There are efficient encoding techniques for arbitrary states under no or only weak constraints [7, 8, 9]. Each of them assumes, however, that an oracle capable of performing an elaborate unitary operation on $\mathcal{O}(2^{n})$ independent states simultaneously is available. This fact prohibits us from finding an efficient encoding scheme for practical use today.
@@
When we tackle a specific problem, however, such techniques for arbitrary states are often found to be too generic; that is, plausible initial states for the problems are characterized by much fewer parameters than $2^{n}.$ Let us consider, for example, a quantum algorithm that admits an initial state in which a probability distribution function is encoded. Even if we want to try various normal distributions as the inputs, a generic encoding technique demands exponential cost, despite the states being specified only by their means and widths. Intuitively, the degrees of freedom in a generic encoding technique are unnecessarily enormous compared to the amount of information specifying an initial state of practical use. Motivated by this consideration, we develop moderately specialized encoding techniques in the present study: generating an arbitrary linear combination (LC) of localized complex functions as a fair trade-off between the degrees of freedom and the computational cost.
@@
To this end, we design the main encoding technique by starting from the probabilistic operation [10, 11] for an LC of discrete Lorentzian functions (LFs), to which we apply the quantum amplitude amplification (QAA) technique [12, 13] to render the encoding deterministic. As is demonstrated later, expanding a target function in LFs is favorable for achieving efficient circuit implementation. Our encoding techniques possess applicability to diverse fields of quantum computation. Amongst them, quantum chemistry in real space [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24] is a promising one, where an initial state can be constructed from one-electron molecular orbitals (MOs) [25, 26], typically expressed as LCs of localized orbitals. We provide resource estimation for such calculations. To demonstrate the efficiency and practical usefulness of our scheme, we perform it on the real quantum computers commercially provided by IBM. Our encoding scheme is for a dense quantum state; that is, it treats a target function that has basically a nonzero amplitude of each computational basis. For encoding schemes for a sparse quantum state, see, e.g., Refs [27, 28, 29] and references therein.
@@
It is noted that encoding techniques exist based on training of parametrized circuits [30, 31]. Although such techniques are promising for practical use, we focus on those based on something other than training in the present study. Also, we note here that Moosa et al.[32] proposed the Fourier series loader method. It encodes first the Fourier coefficients of a target function on a coarse mesh, after which they are reinterpreted as those for a finer mesh to be Fourier transformed backward. See also Ref. [33] for the quantum interpolation method.
@@ II.1.1 Probabilistic operation of an LC of unitaries and amplitude reduction
Figure 1: Each of circuits (a) $\mathcal{C}_{\mathrm{prob}}$ and (b) $\mathcal{C}_{\mathrm{prob}}^{(\mathrm{AR})}$ is for probabilistic operation of an LC $\mathcal{A}$ of unitaries. The explicit construction of the central part $\mathcal{C}$ in $\mathcal{C}_{\mathrm{prob}}$ is described in the original paper [11]. $\mathcal{C}_{\mathrm{prob}}^{(\mathrm{AR})}$ implements the amplitude reduction for the success state by introducing an extra ancilla in addition to those in $\mathcal{C}_{\mathrm{prob}}.$ (c) Each circle represents the weight $W_{\mathrm{f}}$ of failure state in the resultant state of QARA with a QAE error $\Delta w$ as a function of the true weight $w$ of success state in the original state. Each cross represents $W_{\mathrm{f}}$ for QAA with a QAE error. Horizontal lines represent the approximate maximum errors in Eq. (6) for the values of $\Delta w/w.$
@@ II.1.2 Determinization via amplitude reduction and amplification
The generic technique of QAA [12, 13] amplifies the weight of ‘good’ states contained in an input state by rotating it within the two-dimensional subspace spanned by the good states and the orthogonal states, leading to the quadratic speedup for obtaining the good state. This rotation is achieved by the appropriate number of applications of the amplification operator to the input state. In the present study, we adopt this technique for $|\Psi_{\mathrm{AR}}\rangle$ in $\mathcal{C}_{\mathrm{prob}}^{(\mathrm{AR})}$ to raise the success probability. The good states in this case are all those in which the ancillary part is $|0\rangle_{n_{\mathrm{A}}+1}.$ The amplification operator $Q$ for this case can be implemented according to the generic theory. We assume that we already know the weight $w$ of the desired state $|\psi_{\mathrm{lc}}\rangle$ in $\mathcal{C}_{\mathrm{prob}}$ precisely by employing techniques of quantum amplitude estimation (QAE)[13, 34, 35, 36, 37, 38, 39]. Let us consider $m$ applications of $Q$ to $|\Psi_{\mathrm{AR}}\rangle$ by appending the amplification circuit to $\mathcal{C}_{\mathrm{prob}}^{(\mathrm{AR})}.$ By defining $\theta_{w}$ via $w=\sin^{2}\theta_{w}\ (0<\theta_{w}<\pi/2),$ the weight of the good state after the amplification is written as $w_{m}(\theta_{w})\equiv\sin^{2}((2m+1)\theta_{w})$ [12, 13]. If $w_{m}(\theta_{w})$ is equal to unity, the desired state will be obtained with certainty; that is, the nonunitary operation of $\mathcal{A}$ is now deterministic. The operation is still probabilistic otherwise. Specifically, for a fixed $m,$ the following $m$ values of $\theta_{w},$
@@
These observations indicate that $\mathcal{C}_{\mathrm{prob}}^{(\mathrm{AR})}$ with $\theta_{\mathrm{AR,opt}}$ and the subsequent QAA render the nonunitary operation of $\mathcal{A}$ deterministic regardless of $w.$ We refer to this treatment as the determinization of the probabilistic operation and denote it by the quantum amplitude reduction and amplification (QARA) technique. The determinized circuit no longer needs measurements. Since $w$ is on the order of $1/n_{\mathrm{loc}},$ we have $m_{\mathrm{opt}}=\mathcal{O}(\sqrt{n_{\mathrm{loc}}})$ from Eq. (4). In particular, for a case where $1/4\leq w<1,$ we have $m_{\mathrm{opt}}=1.$ Nishi et al. [40] have proposed a similar technique for determinizing the QAA for probabilistic imaginary-time evolution (PITE)[18, 20, 41, 22] steps. The quadratic speedup of PITE [42] with determinization will be a good scheme for nonvariational energy minimization on fault-tolerant quantum computers.
@@ II.1.3 Error analysis for inaccurate QAE
Although we have adopted QARA for the determinization, the weight $w$ of a success state can also be raised to unity by employing the technique in the original paper [13] of QAA, which modifies the reflection circuit at the final iteration. We denote this technique by the modified QAA. Thereby a complex transcendental equation has to be solved using $w$ obtained via QAE to decide the circuit parameters, and the dependence of the error in the amplified state on that in the QAE process is thus difficult to evaluate. Also, the fixed-point search (FPS) technique [43] amplifies the unknown weight $w$ of a success state to $1-\delta^{2}$ for a specified tolerance $\delta.$ $\delta$ must be strictly larger than zero for achieving the quadratic speedup. The modified QAA and FPS techniques do not require an additional qubit in contrast to QARA.
Let us consider a practical situation where the observer obtains an erroneous estimation $w^{(\mathrm{est})}=w+\Delta w$ of the weight $w$ of the original success state. The circuit parameter $\theta_{\mathrm{AR,opt}}$ for amplitude reduction and the repetition number $m_{\mathrm{opt}}$ for amplitude amplification can now involve errors. We demonstrate in Appendix A that QARA allows us to derive a simple relation between the error in the QAE process and that in the determinization. When we approach the limit of $w\to 0$ while keeping $\Delta w/w$ finite, an approximate maximum of the weight of the failure state after the completion of the QARA process is given by
@@
Figure 1(c) shows the weight $W_{\mathrm{f}}$ of the failure state after the QARA process for $\Delta w/w=0.1,0.04,$ and $0.01$ as a function of $w.$ $W_{\mathrm{f}}$ after the QAA process (without amplitude reduction) is also shown in the figure. While the values of $W_{\mathrm{f}}$ from QAA are seen to scatter for each of $\Delta w/w,$ those from QARA exhibit much smaller variations over the wide range of $w.$ Furthermore, $\epsilon_{\mathrm{QARA}}$ approximates $W_{\mathrm{f}}$ well. These observations suggest that QARA is an error-controllable method for determinizing a probabilistic operation.
@@ II.2.2 Displaced single basis function
is easily confirmed to perform modular addition for a computational basis as $T(k)|j\rangle_{n_{q}}=|(j+k)\ \mathrm{mod}\ N\rangle_{n_{q}}.$ The displaced basis function can thus be generated by locating the function at the origin and translating it by $k_{\mathrm{c}\ell}:|\widetilde{f}_{\ell}\rangle=T(k_{\mathrm{c}\ell})U_{\mathrm{orig}}^{(\ell)}|0\rangle_{n_{q}},$ as depicted in Fig. 2(a).
@@ II.2.3 LC of displaced basis functions
with $U_{\mathrm{shift}}^{(\ell)}\equiv U_{\mathrm{shift}}(k_{\mathrm{c}\ell})$ for each $\ell.$ With the LC $\mathcal{L}\equiv\sum_{\ell=0}^{n_{\mathrm{loc}}-1}d_{\ell}\widetilde{U}^{(\ell)}$ of them, the desired state in Eq. (7) is written as $|\psi_{\mathrm{lc}}\rangle=\sum_{\ell=0}^{n_{\mathrm{loc}}-1}d_{\ell}|\widetilde{f}_{\ell}\rangle=\mathrm{QFT}\cdot\mathcal{L}|0\rangle_{n_{q}}.$ $\mathcal{L},$ which is nonunitary in general, admits the probabilistic implementation as a special case of $\mathcal{C}_{\mathrm{prob}}$ in Fig. 1(a). Specifically, the circuit $\mathcal{C}_{\mathrm{lc}}$ in Fig. 2(b) for the data register and
@@
ancillae encodes $|\psi_{\mathrm{lc}}\rangle$ probabilistically. The central part, $\mathcal{C}_{\mathcal{L}},$ implements $\mathcal{L}$ probabilistically by recognizing the $n_{\mathrm{A}}$-bit integer $\ell$ specified by an ancillary state $|\ell\rangle_{n_{\mathrm{A}}}.$ For details, see the original paper [11], which describes the implementation of a generic LC of unitaries. It is noted here that a slight modification to the circuit in the original paper exists in $\mathcal{C}_{\mathcal{L}}.$ That is, since $\mathrm{QFT}^{\dagger}$ in $\widetilde{U}^{(\ell)}$ does not refer to $\ell$ [see Eq. (11)], the controlled $\mathrm{QFT}^{\dagger}$ gates have been replaced by the uncontrolled $\mathrm{QFT}^{\dagger}$ gate, as seen in Fig. 2(b). Since the $\mathrm{QFT}$ operation after $\mathcal{C}_{\mathcal{L}}$ does not act on the ancillae, it is also possible to perform the measurement, observe the success state, and then operate $\mathrm{QFT}$ on the data register.
Let us estimate the depth of $\mathcal{C}_{\mathrm{lc}}$ by assuming that of $U_{\mathrm{orig}}^{(\ell)}$ is independent of $\ell.$ Since $\mathrm{depth}(\mathrm{QFT})=\mathcal{O}(n_{q})$ [45, 46] and $\mathrm{depth}(U_{\mathrm{shift}}^{(\ell)})=\mathcal{O}(1)$ for each $\ell,$ we find
$\mathrm{depth}\left(\mathcal{C}_{\mathrm{lc}}\right)$ $=\mathcal{O}\left(\max\left(n_{\mathrm{loc}}\mathrm{depth}\left(\mathrm{C}^{n_{\mathrm{A}}}U_{\mathrm{orig}}^{(\ell)}\right),n_{q}\right)\right).$ (13)
Although we have assumed the basis functions and the coefficients to be real, it is possible to encode any LC of complex basis functions via complex coefficients by redefining the unitaries and coefficients, as explained in Appendix D. The descriptions below for real values can thus apply to complex values with small modifications for $n_{\mathrm{A}}+1$ ancillae. Also, the protocol described above can be extended to multidimensional space by employing product basis functions, as explained in Appendix E.